Water (3110 g) is heated until it just begins to boil. If the water absorbs 5.01 x 10^5 J of heat in the process. What was the initial temperature of the water ? Express your answer with the appropriate units.


1
Expert's answer
2021-02-09T03:41:32-0500

We use the equation;

Q = m(kg) x C(j/kg) x "\Delta"T(K)


Substituting gives


5.01 x 105J = "\dfrac{3110}{1000}kgx4200J/kg x" "\Delta"T


"\dfrac{5.01 x 10^5J}{3.11 x 4200J}=\Delta"T


"\Delta"T = 38.36K


Assuming that the water begins to boil at 100oC,

Then initial temperature = Final temperature - "\Delta"T


= 100oC - 38.36K = 61.64oC